OpenStudy (anonymous):

Pump A can fill a pool in 12 hours and Pump B in 6 hours. However there is a leak in the pool and a full pool will empty out in 24 hours. How long would it take to fill the pool if pump A and Pump B work and the LEAK WAS STILL PRESENT . Answer pleaseeee .. ASAP Thanks :)

OpenStudy (jack1):

Assume pool = 1000L how many L per hr is pump A pumping? @trix16 ?

OpenStudy (jack1):

...@trix16 ... u there bro?

OpenStudy (anonymous):

There's no liters indicated in the problem just the time

OpenStudy (jack1):

hence "assume"

OpenStudy (jack1):

you can assume it's only a 1L pool if you want or we can just work it out as fractions but easier mental exercise if we assume 1000L and just go through it piece at a time

OpenStudy (jack1):

so assuming a 1000L pool: incoming = Pump A = _______ L/hr + Pump B = _______ L/hr = Total L incoming/hr outgoing = _______L/hr = Total L outgoing/hr Following ok so far dude? @trix16 ?

OpenStudy (jack1):

difference between incoming and outgoing = how fast the pool is filling volume/speed = time

OpenStudy (jack1):

all make sense?

OpenStudy (anonymous):

So what's the answer?

OpenStudy (jack1):

u tell me and i'll tell u if you're right i gave you the solution above...

OpenStudy (jack1):

the goal here is for you to understand, so you can work these type of questions out, ...not to give you the answer to a specific question and wait so we can then give you the answer to your next question

OpenStudy (anonymous):

Wait, i will solve for it

OpenStudy (anonymous):

4 Hours @Jack1 ?

OpenStudy (jack1):

close, but no, how did u arrive at that?

OpenStudy (anonymous):

Wait, 20 hours?

OpenStudy (jack1):

no

OpenStudy (jack1):

assume 1000L Pump A can pump 1000L in 12 hrs 1000L/12 hrs = 83.33L/hr Pump B can pump 1000L in 6 hrs 1000L/6 hrs = 166.6L/hr

OpenStudy (jack1):

pool is leaking at 1000L every 24hrs 1000L / 24hrs = 41.66L/hr

OpenStudy (jack1):

so total outgoing = 41.66L/hr total incoming = 83.33+166.66 = 250L/hr

OpenStudy (jack1):

difference between incoming and outgoing = how fast the pool is filling incoming - outgoing = speed pool is filling 250 - 41.66 = 208.33L/hr = how fast the pool is filling

OpenStudy (jack1):

volume/speed = time Volume of pool/ speed of overall flow = time to fill pool 1000L / 208.33 = ___hrs = answer

OpenStudy (jack1):

can u get it from here dude?

OpenStudy (anonymous):

Yes, yes :)

OpenStudy (jack1):

sweet, seeya

OpenStudy (anonymous):

Thanks :) @jack1

OpenStudy (jack1):

all good dude @trix16 ;)